OpenStudy (anonymous):

The distance between two cities is ninety miles, and a woman drives from one city to the other at a rate of 45 mph. At what rate must she return if the total travel time is three hours and forty minutes?

OpenStudy (agent47):

90/45mph=x(h)

OpenStudy (agent47):

90/v=x+x1

OpenStudy (agent47):

x+x1=225 mins

OpenStudy (agent47):

Time to get to the city was 2 hours. So she has to get back in 1 hours and 40 mins

OpenStudy (anonymous):

it took her \(\frac{90}{45}=2\) hours to get there. she has only 1 hour and 40 minutes or \(1\tfrac{2}{3}\) hours left to make the return trip. to get the rate, divide the distance by the time. i.e divide \[90\div 1\tfrac{2}{3}\]

OpenStudy (agent47):

well I'll finish what I was saying, though someone got there first

OpenStudy (anonymous):

\[90\div 1\tfrac{2}{3}=90\div \frac{5}{3}=90\times \frac{3}{5}\]

OpenStudy (agent47):

ok I won't finish what I was saying lol